# MMBench-Video: A Long-Form Multi-Shot Benchmark for Holistic Video Understanding

### v1

@@ Abstract
{NoHyper}footnotetext: These authors contributed equally to this work.{NoHyper}footnotetext: The work was conducted during an internship at Shanghai AI Laboratory.{NoHyper}footnotetext: Corresponding author.
@@ 1 Introduction
To address these problems, we develop a new VideoQA benchmark, MMBench-Video, to evaluate the effectiveness of LVLMs in video understanding. It incorporates approximately 600 web videos with rich context from YouTube, spanning 16 major categories, including News, Sports, etc., covering most video topics people watch in their daily lives. Each video ranges in duration from 30 secs to 6 mins, to accommodate the evaluation of video understanding capabilities on longer videos. The benchmark includes roughly 2,000 original question-answer (QA) pairs, contributed by volunteers, covering a total of 26 fine-grained capabilities. During dataset collection, we implement quality control strategies to explicitly increase the proportion of temporal indispensable questions-1-1-1A visual question is temporal indispensable if it can not be correctly solved by viewing any random frame. . Quantitative statistics show that MMBench-Video significantly differs from existing benchmarks in terms of temporal duration, context richness, and temporal indispensability.
@@ 2.2 Video Question Answering
Video Question Answering (VideoQA) is a critical method for assessing the depth of understanding that models possess regarding video content. The research community has progressively developed a wide array of VideoQA benchmarks, spanning various visual domains such as movies [46], TV shows [28, 21], video games [37], synthetic scenarios [57], and egocentric videos [20]. These benchmarks typically assess models trained on their respective training sets, demanding concise answers for evaluation. However, Large Vision and Language Models (LVLMs), which are often not trained on domain-specific data, face challenges in adapting to these benchmarks due to their diverse answer styles. To mitigate this, Video-ChatGPT [36] employs GPT-3.5 as a scoring mechanism for free-form responses from VLMs. The method was applied to evaluate several popular benchmarks [58, 24, 52], covering topics including concept existence, objection relationship, and activity recognition. Despite its broad adoption [32, 54, 45], this approach is limited by suboptimal accuracy and stability, as well as poor alignment with human preferences. Additionally, those benchmarks primarily consist of short videos, which contrasts with the typical length of web videos. In response, we present MMBench-Video, a novel dataset tailored for longer videos, challenging models to generate detailed, free-form responses to complex questions. We adopt GPT-4-based evaluation, which improves correctness and robustness, offering a more stable evaluation strategy compared to previous methods.
@@ 3.2 Dataset Statistics
Benchmark MSVD TGIF MSRVTT ActivityNet MMBench-Video
Input Frames 1 8 1 8 1 8 1 8 1 8
Original Score 2.62 2.93 2.66 3.18 2.01 2.33 2.65 3.05 0.78 1.63
Normalized Score 52.4 58.6 53.2 63.6 40.2 46.6 53.0 61.0 26.0 54.3
@@
Table 2: Comparing the temporal indispensability of existing VideoQA benchmarks. MMBench-Video adopts a different grading paradigm compared to other benchmarks (3-grade vs. 5-grade). We calculate the ‘Normalized Score’ (normalize to 0-100) to ease the cross-benchmark comparisons. Benchmarks with smaller 1-frame scores and Score-[1f] / Score-[8f] ratio feature better temporal indispensability.
@@
To quantitatively measure the temporal indispensability of each VideoQA benchmark, we randomly sample 1000 QAs from MSVD, TGIF, MSRVTT, and ActivityNet, and conduct a study on the subsets as well as MMBench-Video. We evaluate GPT-4o (by far the most powerful LVLM) on these benchmarks under 1-frame and 8-frame settings, and present the results in Tab. 2. Notably, GPT-4o using a 1-frame input achieves a normalized score of approximately 50%, retaining over 80% of its performance compared to an 8-frame input across MSVD, TGIF, MSRVTT, and ActivityNet. In contrast, when assessed on the MMBench-Video, GPT-4o using a 1-frame input preserves only 47.8% of its efficacy compared to its performance with an 8-frame input, yielding a normalized score of just 26.0%. This marked difference underscores the temporal importance of MMBench-Video.
@@ 4 Experiment
Table 3: Evaluation Result of Video Models on MMBench-Video. CP, FP[S], FP[C], HL stands for four L-2 perception capabilities: Coarse Perception, Single-Instance Finegrained Perception, Cross-Instance Finegrained Perception, and Hallucination. LR, AR, RR, CSR, TR stand for five reasoning capabilities: Logic, Attribute, Relation, Commonsense, and Temporal Reasoning. All scores are based on a 3-grade marking scheme: 0 for worst, 3 for best. -[$N$f] indicates the method take $N$ frames uniformly sampled from a video as input. Among Open-Source Video-LLMs, ShareGPT4Video-8B-[16f*] follows the IG-VLM [26] strategy and presents 16 frames in a $4\times 4$ grid. VideoStreaming-[64f+] accepts streaming videos and takes at least 64 frames as inputs.
@@ 4.1 Main Results
Open-Source Video-LLMs. We first identify and evaluate representative open-source Video-LLMs using MMBench-Video. Adhering to their default settings, these Video-LLMs process a sequence of video frames, with the number of frames varying from eight to dozens. Interestingly, we observe that all Video-LLMs exhibit comparably subpar performance on MMBench-Video, despite notable performance disparities on other benchmarks. For instance, VideoChat2 surpasses Video-ChatGPT by 18% on the MSVD-QA score (3.9 vs. 3.3), yet the performance gap narrows to just 6% on the MMBench-Video score (0.99 vs. 0.93). All video LLMs attain an average score close to 1 (out of a total of 3), with the top-performing model VideoStreaming [43] reaching a mere 1.12. These findings suggest that the current state of video models’ proficiency in understanding MMBench-Video is nascent, underscoring the challenges and emphasizing the necessity for advancements in video LLMs to enhance their capability and effectiveness in interpreting varied video content.
@@
Open-Source LVLMs for Images. A significant number of LVLMs [17, 34, 5, 55, 27, 12, 41, 50] have been developed to comprehend image content and execute visual reasoning tasks. During our evaluation, we focused on LVLMs that support the multi-image inference interface. We assess four prominent open-source LVLMs: Idefics2-8B [27], Qwen-VL-Chat [5], mPLUG-Owl2 [56], and InternVL-Chat-v1.5 [12], using MMBench-Video. To ascertain the models’ ability to effectively leverage multiple input frames, we evaluated them under two distinct settings: 1-frame and 8-frame inputs. Results in Tab. 3 indicate that all models, except Qwen-VL-Chat, exhibit a substantial enhancement in performance when utilizing 8 frames compared to a single frame. Notably, InternVL-Chat-v1.5 emerges as the top performer, achieving an impressive average score of 1.26 with 8 frames as inputs, significantly surpassing all other evaluated video LLMs.
@@
Proprietary LVLMs for Images. Unlike their open-source counterparts, most proprietary LVLMs accept arbitrary interleaved images and text as input. However, the context window size still limits the maximum number of images. We evaluate several proprietary LVLMs, including Claude-3v, Gemini-Pro-v[1.0/1.5], GPT-4v, and GPT-4o on MMBench-Video with varying numbers of frames. Claude-3v struggles with 8-frame inputs and is only evaluated under the 4-frame setting. As anticipated, it exhibits the poorest performance among proprietary LVLMs when handling multiple frames. In contrast, other proprietary models demonstrate notably superior performance compared to the state-of-the-art open-source InternVL-Chat. Particularly impressive is GPT-4o, which, when processing 16 frames, achieves an outstanding overall score of 1.86. This result positioned GPT-4o 66% ahead of the best open-source video LLM and 48% ahead of the best open-source image LVLM.
@@ 4.3 Incorporating Speech Further Improves Proprietary LVLMs
(a) Results of Video-LLaVA.
(b) Results of GPT-4o.
Figure 6: Rating distribution of different judge models. We report the results of Video-LLaVA and GPT-4o respectively judged by GPT-3.5-Turbo (1106) and GPT-4-Turbo (1106).
@@ 4.4 The Superior Performance of GPT-4 as a Judge
Due to the discrepancy between the predictions of Video-LLM and the ground truth answers, existing VideoQA benchmarks largely rely on a judge model to assess the model responses. The capability of judge models can significantly influence the final results. To quantitatively study the impact of judge models, we utilize different versions of GPT-3.5 and GPT-4 for evaluation and report the results in Tab. 7 and Fig. 6. We observe that GPT-3.5 tends to assign high scores (typically 2 and 3), potentially leading to inaccuracies. To investigate the alignment with human preferences across different judge models, we conduct study based on a randomly selected subset of 100 questions. Two of the authors manually rate the responses from Video-LLaVA and GPT-4o, and we then report the mean absolute error between different judge models and the averaged human ratings. Tab. 7 shows that GPT-3.5 exhibits a significantly larger discrepancy with human preferences and greater inter-version variance.
@@ 5 Conclusion
Appendix A Additional Experiments
A.1 The Impact of Incorporating Speech over the entire MMBench-Video
In the main paper, we report the quantitative results of speech improvement on the subset of videos with subtitles available from YouTube. In MMBench-Video, approximately half of videos do not include parseable video title tracks. In Tab. 8, we report the impact of incorporating speech information across the entire MMBench-Video dataset. While only half of the VideoQAs are enhanced with speech, the overall performance improvement on the full benchmark remains significant. It is evident that the enhancement in reasoning capabilities surpasses that of perceptual abilities. Speech typically conveys contextual information absent in static visual inputs, facilitating further reasoning by the VLM. Meanwhile, improvements in coarse perception are minimal or remain largely unchanged (for GPT-4o-[16f]). This can be attributable to the fact that perception is predominantly reliant on visual inputs, and speech information does not significantly augment the model’s performance in coarse perception.
@@ A.1 The Impact of Incorporating Speech over the entire MMBench-Video
A.2 Detailed Analysis of L-2 Capability
@@
Appendix B Additional Dataset Analysis
@@
B.1 Temporal Dispensable Data Filtering with LVLM
@@
B.2 Question Type Analysis
@@
Figure 8: Visualization of Samples in MMBench-Video. Part 1 out of 3.
Figure 9: Visualization of Samples in MMBench-Video. Part 2 out of 3.
Figure 10: Visualization of Samples in MMBench-Video. Part 3 out of 3.
Appendix C Prompts Adopted in MMBench-Video
@@
C.1 System Prompt for GPT-based Evaluation
@@
C.2 System Prompt for the Inference of LVLMs with Multi-Frame Inputs
@@
C.3 Input Prompt Template for the Inference of LVLMs with Multi-Frame Inputs
@@
Appendix D Limitations and Broader Impacts

### v2

@@ 1 Introduction
To address these problems, we develop a new VideoQA benchmark, MMBench-Video, to evaluate the effectiveness of LVLMs in video understanding. It incorporates approximately 600 web videos with rich context from YouTube, spanning 16 major categories, including News, Sports, etc., covering most video topics people watch in their daily lives. Each video ranges in duration from 30 secs to 6 mins, to accommodate the evaluation of video understanding capabilities on longer videos. The benchmark includes roughly 2,000 original question-answer (QA) pairs, contributed by volunteers, covering a total of 26 fine-grained capabilities. During dataset collection, we implement quality control strategies to explicitly increase the proportion of temporal indispensable questions [^1] A visual question is temporal indispensable if it can not be correctly solved by viewing any random frame. . Quantitative statistics show that MMBench-Video significantly differs from existing benchmarks in terms of temporal duration, context richness, and temporal indispensability.
@@ 2.2 Video Question Answering
Video Question Answering (VideoQA) is a critical method for assessing the depth of understanding that models possess regarding video content. The research community has progressively developed a wide array of VideoQA benchmarks, spanning various visual domains such as movies [50], TV shows [29, 22], video games [40], synthetic scenarios [61], and egocentric videos [21]. These benchmarks typically assess models trained on their respective training sets, demanding concise answers for evaluation. However, Large Vision and Language Models (LVLMs), which are often not trained on domain-specific data, face challenges in adapting to these benchmarks due to their diverse answer styles. To mitigate this, Video-ChatGPT [39] employs GPT-3.5 as a scoring mechanism for free-form responses from VLMs. The method was applied to evaluate several popular benchmarks [62, 25, 56], covering topics including concept existence, objection relationship, and activity recognition. Despite its broad adoption [34, 58, 48, 65], this approach is limited by suboptimal accuracy and stability, as well as poor alignment with human preferences. Additionally, those benchmarks primarily consist of short videos, which contrasts with the typical length of web videos. In response, we present MMBench-Video, a novel dataset tailored for longer videos, challenging models to generate detailed, free-form responses to complex questions. We adopt GPT-4-based evaluation, which improves correctness and robustness, offering a more stable evaluation strategy compared to previous methods.
@@ 3.2 Dataset Statistics
      Benchmark       MSVD       TGIF       MSRVTT       ActivityNet
      Input Frames       1       8       1       8       1       8       1       8
      Original Score       2.62       2.93       2.66       3.18       2.01       2.33       2.65       3.05
      Normalized Score       52.4       58.6       53.2       63.6       40.2       46.6       53.0       61.0
@@
      Benchmark       EgoSchema       Video-MME*       Next-GQA       MMBench-Video
      Input Frames       1       8       1       8       1       8       1       8
      Original Score       0.65       0.70       0.54       0.68       0.78       0.84       0.78       1.63
      Normalized Score       65.0       70.0       54.0       68.0       78.0       84.0       26.0       54.3
[…]
Table 2: Comparing the temporal indispensability of existing VideoQA benchmarks. MMBench-Video adopts a different grading paradigm compared to other benchmarks (3-grade vs. 5-grade). We calculate the ‘Normalized Score’ (normalize to 0-100) to ease the cross-benchmark comparisons. Benchmarks with smaller 1-frame scores and Score-[1f] / Score-[8f] ratio feature better temporal indispensability. For Video-MME, we excluded long-duration videos in this comparison to better align with other datasets.
@@
To quantitatively measure the temporal indispensability of each VideoQA benchmark, we randomly sample 1000 QAs from orginal VideoQA datasets and comprehensive video understanding benchmark, and conduct a study on the subsets as well as MMBench-Video. We evaluate GPT-4o (by far the most powerful LVLM) on these benchmarks under 1-frame and 8-frame settings, and present the results in Tab. 2. Notably, GPT-4o using a 1-frame input achieves a normalized score of approximately 50%, retaining over 75% of its performance compared to an 8-frame input across previous VideoQA datasets. Even the latest benchmarks for comprehensive video understanding struggle to fully assess a model’s temporal capabilities. In contrast, when assessed on the MMBench-Video, GPT-4o using a 1-frame input preserves only 47.8% of its efficacy compared to its performance with an 8-frame input, yielding a normalized score of just 26.0%. This marked difference underscores the temporal importance of MMBench-Video.
@@ 4 Experiment
GPT-4o-[1fps] 2.15 2.23 2.24 2.01 1.90 2.19 2.11 2.12 2.17 1.94 1.97 2.08
Table 3: Evaluation Result of Video Models on MMBench-Video. CP, FP[S], FP[C], HL stands for four L-2 perception capabilities: Coarse Perception, Single-Instance Finegrained Perception, Cross-Instance Finegrained Perception, and Hallucination. LR, AR, RR, CSR, TR stand for five reasoning capabilities: Logic, Attribute, Relation, Commonsense, and Temporal Reasoning. All scores are based on a 3-grade marking scheme: 0 for worst, 3 for best. -[$N$f] indicates the method take $N$ frames uniformly sampled from a video as input. -[$N$fps] indicates the method uses $N$ frames per second uniformly sampled from a video as input. Among Open-Source Video-LLMs, ShareGPT4Video-8B-[16f*] follows the IG-VLM [27] strategy and presents 16 frames in a $4\times 4$ grid. VideoStreaming-[64f+] accepts streaming videos and takes at least 64 frames as inputs.
@@ 4.1 Main Results
Open-Source Video-LLMs. We first identify and evaluate representative open-source Video-LLMs using MMBench-Video. Adhering to their default settings, these Video-LLMs process a sequence of video frames, with the number of frames varying from eight to dozens. Interestingly, we observe that all Video-LLMs exhibit comparably subpar performance on MMBench-Video, despite notable performance disparities on other benchmarks. For instance, VideoChat2 surpasses Video-ChatGPT by 18% on the MSVD-QA score (3.9 vs. 3.3), yet the performance gap narrows to just 6% on the MMBench-Video score (0.99 vs. 0.93). All video LLMs attain an average score close to 1 (out of a total of 3), with the top-performing model LLaVA-NeXT-Video [64] reaching a mere 1.14. These findings suggest that the current state of video models’ proficiency in understanding MMBench-Video is nascent, underscoring the challenges and emphasizing the necessity for advancements in video LLMs to enhance their capability and effectiveness in interpreting varied video content.
@@
Open-Source LVLMs for Images. A significant number of LVLMs [18, 37, 6, 59, 28, 13, 44, 54] have been developed to comprehend image content and execute visual reasoning tasks. During our evaluation, we focused on LVLMs that support the multi-image inference interface. We assess several prominent open-source LVLMs: Idefics2-8B [28], Qwen-VL-Chat [6], mPLUG-Owl2 [60], InternVL-Chat-v1.5 [13], InternVL2 [13] and VILA1.5 [35] using MMBench-Video. To ascertain the models’ ability to effectively leverage multiple input frames, we evaluated them under two distinct settings: 1-frame and 8-frame inputs. Results in Tab. 3 indicate that all models, except Qwen-VL-Chat, exhibit a substantial enhancement in performance when utilizing 8 frames compared to a single frame. Notably, VILA-1.5-40B emerges as the top performer, achieving an impressive average score of 1.61 with 14 frames as inputs, significantly surpassing all other evaluated video LLMs.
@@
Proprietary LVLMs for Images. Unlike their open-source counterparts, most proprietary LVLMs accept arbitrary interleaved images and text as input. We evaluate several proprietary LVLMs, including Claude-3v, Gemini-Pro-v[1.0/1.5], GPT-4v, and GPT-4o on MMBench-Video with varying numbers of frames. Claude-3v struggles with 8-frame inputs and is only evaluated under the 4-frame setting. As anticipated, it exhibits the poorest performance among proprietary LVLMs when handling multiple frames. In contrast, other proprietary models demonstrate notably superior performance compared to the state-of-the-art open-source VILA-1.5. Particularly impressive is GPT-4o, which, when processing 16 frames, achieves an outstanding overall score of 1.86. This result positioned GPT-4o 63% ahead of the best open-source video LLM and 16% ahead of the best open-source image LVLM. To assess the potential of advanced proprietary models, we experiment with the fps sampling method. By increasing the number of video frames, the model’s perception improves as adjacent frames provide mutual support, resulting in more accurate interpretations. This approach also captures previously missed content, enhancing the model’s ability to answer questions reliant on such information and boosting its reasoning skills. The only observed drawback is a slight increase in hallucination, possibly due to excessive video content leading to responses not grounded in reality.
@@ 4.4 The Superior Performance of GPT-4 as a Judge
Due to the discrepancy between the predictions of Video-LLM and the ground truth answers, existing VideoQA benchmarks largely rely on a judge model to assess the model responses. The capability of judge models can significantly influence the final results. To quantitatively study the impact of judge models, we utilize different versions of GPT-3.5 and GPT-4 for evaluation and report the results in Tab. 7. We observe that GPT-3.5 tends to assign higher scores (typically 2 and 3), which can result in inflated final results and potential inaccuracies. To investigate the alignment with human preferences across different judge models, we conduct study based on a randomly selected subset of 100 questions. Two of the authors manually rate the responses from Video-LLaVA and GPT-4o, and we then report the mean absolute error between different judge models and the averaged human ratings. Tab. 7 shows that GPT-3.5 exhibits a significantly larger discrepancy with human preferences and greater inter-version variance. In contrast, Qwen2-72B-Instruct aligned more closely with human ratings, suggesting the potential of using advanced open-source LLMs as evaluators. Compared to the other two, GPT-4 demonstrated greater resistance to manipulation and fairly evaluated predictions without bias, supporting its role as an evaluator for the MMBench-Video benchmark.
@@ 5 Conclusion
Acknowledgement
This project is supported by the National Key R&D Program of China (No.2022ZD0161600), the Shanghai Postdoctoral Excellence Program (No.2023023), China Postdoctoral Science Fund (No.2024M751559), and Shanghai Artificial intelligence Laboratory.
Appendix A OpenSource Datasets and Codes
We have uploaded the full MMBench-Video dataset to HuggingFace, you can access the data at https://huggingface.co/datasets/opencompass/MMBench-Video. The code related to performance evaluation of Video-LLM and LVLM using MMBench-Video has been released in VLMEvalKit and the test results have been published in OpenVLM Video Leaderboard. We provide the DataSheet at the end of this document.
Author Statement and Data Licence. The authors bear all responsibility in case of violation of rights and confirm that this dataset is opensourced under the Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license.
Appendix B Additional Experiments
B.1 The Impact of Incorporating Speech over the entire MMBench-Video
In the main paper, we report the quantitative results of speech improvement on the subset of videos with subtitles available from YouTube. In MMBench-Video, approximately half of videos do not include parseable video title tracks. In Tab. 8, we report the impact of incorporating speech information across the entire MMBench-Video dataset. Initially, we tested GPT-4o without visual input, and the results showed that even advanced models struggle to solve most problems using only question inputs. However, with subtitles, the model’s ability to visualize content improves, boosting performance by 200%. Further experiments with 8 and 16 frame inputs show that, even though only half of the VideoQAs include speech, the overall performance on the full benchmark remains significantly enhanced. While only half of the VideoQAs are enhanced with speech, the overall performance improvement on the full benchmark remains significant. It is evident that the enhancement in reasoning capabilities surpasses that of perceptual abilities. Speech typically conveys contextual information absent in static visual inputs, facilitating further reasoning by the VLM. Meanwhile, improvements in coarse perception are minimal or remain largely unchanged (for GPT-4o-[16f]). This can be attributable to the fact that perception is predominantly reliant on visual inputs, and speech information does not significantly augment the model’s performance in coarse perception.
@@ B.1 The Impact of Incorporating Speech over the entire MMBench-Video
B.2 Detailed Analysis of L-2 Capability
@@
B.3 Model Performance across Video Time Length and Video Shots
The length of the video and the number of shots are indeed key factors affecting model performance. Videos with fewer and shorter shots may perform better at lower frame counts, while longer or multi-shot videos require more visual content to fully leverage the model’s capabilities.
Fig. 6 illustrates the trend of model scores in relation to shot count and video length. It is evident that the performance of GPT-4o, sampled at different frame rates, declines as video length increases, whereas the performance of open-source models such as InternVL-Chat-v1.5 and Video-LLaVA remains relatively stable. Compared to video length, model performance is more significantly influenced by the number of video shots. With over 50 shots in a video, the performance of GPT-4o drops to 75% of its original score. This suggests that the model’s performance is more closely tied to the number of shots than to video length, as the frequent shot transitions make it more challenging for the model to comprehend the video, resulting in lower scores.
Figure 6: Performance across varying video lengths and video shots. We report the results of InternVL-Chat-V1.5-[1/8f], GPT-4o-[1/8f] and Video-LLaVA judged by GPT-4-Turbo (1106).
Appendix C Additional Dataset Analysis
@@
C.1 Temporal Dispensable Data Filtering with LVLM
@@
C.2 Question Type Analysis
@@
Appendix D Prompts Adopted in MMBench-Video
@@
System Prompt for GPT-based Evaluation.
@@
System Prompt for the Inference of LVLMs with Multi-Frame Inputs.
@@
Input Prompt Template for the Inference of LVLMs with Multi-Frame Inputs.
@@
Appendix E Limitations and Broader Impacts